OpenStudy (anonymous):

The speed of sound in methanol is 1135 m/s. What is the bulk modulus of methanol?

OpenStudy (anonymous):

v=\[\sqrt{k/ density}\] 1288225\[\times\]density of methanol 1288225 \[\times\] 791.30 =1019372443

OpenStudy (anonymous):

Thank-you!